a) Use the Frobenius method to obtain the general solution of the following equation, valid near x=0.

b)

Exercise 4.4.7 Determine the two values of the constant a for which all solutions of can be written as v(x) = >A4r* k-0
Answer #1

х (2-1) tthe trobenius AUies abad -0is nt n 0 n+Y- ntr)Cn 2) n=0 (ntYn+Y-) c, n-2 (a) n=0 values (n(n-)C Y- (x-)(n ) n= 0 n(nt)(ntY-) ntY- (ntr) n+Y х n n=0 n o κΣι,χ n+ 0 _ indiaial quation conaida the last paver Aum Hesc Aum inJeaat poe Aum n= 04ndicial equaton (7-) Y= 0 Y- 2 > 0 Value of athis fisat term t becomes 0 um non n in rom M+(nt- )- (nt)Co x- n= n+r)-C x =becomes Y n+(n+¥ -1) - (n+l+x)|C nt) n+Y n+x Y+)Cn+ n+Y = 0 + n 0 nY ntr)-Cn+Y 0 nt) Scanned with CS CamScannerd-(n+) G |Cn1 (ntrt) (n+Y)-(n+Y 01,2 - (n)n nro+(no)-(nto d-n) Cn nn-(n+) Jo - (n+)Cn Cnti n+A1(n+d) - (n+2+) o-(n+2)Cr ntl nvulne к. calonlate Asgncun Cs we can thae Cm nthing and n+y we CS Scanned with CamScanner